Question 3 1 pts In order for your clients to reach their retirement goal, lets say youve estimated that they need to have
Answer #1

Inflation Adjusted Return= ((1+Return)/ (1 +Inflation Rate))-1

Inflation Adjusted Return=((1+0.067)/(1+0.034))-1

=(1.067/1.034)-1

=1.031915

Client has already saved 1,30,000$.

Future value of this after 26 years= 1,30,000*(1+0.031915)^26

=2,94,228$

Amount of future value to be covered= 28,00,000-2,94,228

=25,05,772$

Annual savings= Future value/((1+inflation adjusted return)^no of yrs)

=25,05,772/(1.031915)^26

=25,05,772/17.4892

=1,43,275$

Amount to be saved yearly= 1,43,275$